2. The 52 point totals for the previous superbowls are shown below. x = 2434> χ2 = 123,274 stem-and-leaf n = 49 Leaf Unit 1.0 1. Calculate mean and standard deviation Mean= 2 123 2 79 3 01114 3 67778899 4 1344 4 55667778 5 0012234 5 56669 6 112 6 55699 2. What is the sample range of the data? 3. Calculate a trimmed mean and denote what percent it is? Trimmed Mean Percent ? Below.what point total would be a mild outlier?
Answer #1

1)

mean = sum x /n = 2434/52 = 46.8076923

2

S^2= (123274 - 52 * 46.8076923 ^2 )/ 51

= 183.2172

S = 13.53577

2)

range = maximum -minimum

= 75 - 21

=54